Suppose a population is composed of only 3 measures: 1, 2, and 3. The possible samples of size 2 can be draw from this population. List all the possible sample size 2 when repetition is allowed or with replacement. With solution

1
Expert's answer
2022-03-20T06:41:53-0400

All the samples are:

"\left( 1,1 \right) ,\left( 1,2 \right) ,\left( 1,3 \right) \\\left( 2,1 \right) ,\left( 2,2 \right) ,\left( 2,3 \right) \\\left( 3,1 \right) ,\left( 3,2 \right) ,\left( 3,3 \right)"